THE ROTO-TORSIONAL SPECTRUM OF GLYCOLALDEHYDE

\underline{M.L. SENENT} , H.MASSO , Departamento de Astrofisica Molecular e Infrarroja, Instituto de Estructura de la Materia, C.S.I.C., Madrid 28006, Spain ; M.VILLA , Departamento de Quimica, UAM-I Purisima y Michoacan, s/n , CP 09340, Mexico D.F ; R. DOMINGUEZ-GOMEZ , Departamento de Ingenieria Civil, Catedra de Quimica, E.U.I.T. Obras Publicas, Universidad Politecnica de Madrid.

Glycolaldehyde is the first interstellar monosaccharide, that has been detected in emission toward the Galactic center source Sgr B2(N) . The molecule presents four conformers, three show C_{s} symmetry (I, II and IV) and one, III, represents a double minimum of C_{1} symmetry. Two minima I and IV are cis forms whereas II and III are trans-forms. Cis-I is the most stable geometry that stabilizes by the formation of a hydrogen bond between the OH and the C=O carbonyl group. The relative energies of the remaining minima have been determine to be 1278.2 cm^{−1} (II), 1297.2 cm^{−1} (III) and 1865.2 cm^{−1} (IV) using MP4(SDTQ)/cc-pVQZ. The cis-trans transformation is hindered by a barrier of 1895 cm^{−1} and carries out a significative change of the rotational constants. The roto-torsional energies are obtained by solving the roto-torsional Hamiltonian depending on two coordinates. The potential energy surface is calculated from the MP4/cc-pVQZ energies of a grid of 74 conformations. The rotational constants and the distortion centrifugal constants are obtained in the zero point torsional energies of the cis and the trans forms.